Two closely wound circular coils have the same number of turns, but one has twice the

radius of the other. How are the self-inductances of the two coils related? Explain your

reasoning.



1
Expert's answer
2022-03-11T08:33:39-0500

Answer

Self inductance of cool is given by

"L=\frac{\mu_0N^2\pi r^2}{l}"

Now of first coil

"L_1=\frac{\mu_0N^2\pi r_1^2}{l}"

Now for second coil

"r_2=2r_1"

So "L_2=\frac{\mu_0N^2\pi r_2^2}{l}\\"


"L_2=4L_1"

This is relationship between both coil.